need to know the slope and y intercept of these two problems
1. y=-4 -x
2. x+y=12
please email me asap or at your earliest time my email is cstayfly2.cr@gmail.com

Answer by rfer(16322)

1)
slope m=-1
y int b=-4
(0,-4)
2)
y=-x+12
m=-1
b=12
(0,12)